# Spectrally Deconfounded Random Forests

### v2

@@ Abstract
We introduce a modification of Random Forests to estimate functions when unobserved confounding variables are present. The technique is tailored for high-dimensional settings with many observed covariates. We employ spectral deconfounding techniques to minimize a deconfounded version of the least squares objective, resulting in the Spectrally Deconfounded Random Forests (SDForests). We show how the omitted variable bias in estimating a direct effect goes to zero, assuming dense confounding and high-dimensional data. We compare the performance of SDForests with that of classical Random Forests in a simulation study and a semi-synthetic setting using single-cell gene expression data. Empirical results suggest that SDForests outperform classical Random Forests in estimating the direct regression function, even if the theoretical assumptions are not perfectly met, and that SDForests and classical Random Forests have comparable performance in the non-confounded case. We provide an R-Package for SDForest, and supplementary materials for this article are available online.
@@ 1 Introduction
Random Forests (Breiman 2001) and their variations, such as Random Survival Forests (Hothorn 2005, Taylor 2011), Quantile Regression Forests (Meinshausen 2006), or distributional versions of Random Forests (Hothorn & Zeileis 2021, Ćevid et al. 2022) are successfully applied to a wide range of datasets. In many cases of observational data, however, problems with “omitted variable bias” (Cinelli & Hazlett 2020, Wilms et al. 2021) occur. This means that a bias is induced in estimating relationships using standard Random Forest versions when covariates that correlate with other covariates and the response are not observed and included. In the setting of causality, this can be viewed as a confounded causal relationship with unobserved confounders (Pearl 2009, Peters et al. 2016). A popular approach to deal with unobserved confounding is to use instrumental variables (IV) regression techniques (Bowden et al. 1990, Angrist et al. 1996, Stock & Trebbi 2003). Finding enough strong and valid instrumental variables can be challenging, especially if many covariates with potential effects on the response are observed since the number of instruments must be as large as the number of effective covariates. Another possible way of reducing the hidden confounding bias, which we will adopt here, is to make some kind of “dense confounding effect” assumption, meaning that the non-observed factors or confounders affect most of the covariates. Then, a standard approach is to estimate the hidden confounding using methods from high-dimensional factor analysis (Bai 2003) and explicitly adjust for them, see for example Leek & Storey (2007), Gagnon-Bartsch & Speed (2012), Fan et al. (2024) for approaches in this direction. Instead of estimating the latent factors explicitly, one can adjust for them implicitly using spectral transformations (Ćevid et al. 2020). Applying such spectral transformations, especially the trim transform, which we introduce later, does not require any tuning and is computationally very fast since it is a simple and explicit function of the singular value decomposition of the design matrix.
@@ 1.1 Our Contribution
Our contribution is a Random Forest algorithm that is able to address, at least partially, the problem of hidden confounding. Our proposal combines the great advantage and flexibility of standard Random Forests with spectral deconfounding techniques for dealing with the bias from unobserved factors or confounding. The latter was originally proposed for linear models by Ćevid et al. (2020), Guo et al. (2022) and was further developed for nonlinear additive models by Scheidegger et al. (2025). The application of spectral deconfounding techniques for Random Forests is novel.
We develop a new algorithm with R-package SDModels (Ulmer & Scheidegger 2025) and show its performance in estimating the direct and unconfounded relationship between the observed covariates and the response. We demonstrate that in presence of hidden confounding, our method, the Spectrally Deconfounded Random Forest (SDForest), outperforms the standard Random Forests in many aspects. If no latent factor or confounding exists, our SDForests and standard Random Forests perform similarly, perhaps with a minimal edge in favor of the classical algorithm. Thus, if one is unsure whether hidden confounding is present, there is much to gain but not much to lose.
@@ 2 Confounding Model
Here, $X\in\mathbb{R}^{p}$ are the predictors, $Y\in\mathbb{R}$ is the response and $H\in\mathbb{R}^{q}$ are unobserved hidden confounding factors. We assume that the confounder influences $X$ with a linear effect $\Gamma\in\mathbb{R}^{q\times p}$ and $Y$ with a linear effect $\delta\in\mathbb{R}^{q}$ (see Appendix C for a note on non-linear confounding); without loss of generality, we can assume $\mathrm{Cov}(H)=I$. Furthermore, $\nu$ is a random variable with mean zero and variance $\sigma_{\nu}^{2}$, and $E$ is a random vector with mean zero and covariance $\Sigma_{E}$, and $E$ and $H$ are uncorrelated. The error term $E$ can be viewed as the unconfounded predictor if $\Gamma$ equals zero. Finally, $f^{0}\in\mathcal{F}$, where $\mathcal{F}$ is some class of functions from $\mathbb{R}^{p}$ to $\mathbb{R}$. The function $f^{0}$ encodes the direct causal relationship of interest, describing the causal relation of $X$ on $Y$. The described model is visualized as a graph in Figure 1.
$X$$Y$$H$$f^{0}$$\Gamma$$\delta$
@@ 3 Generic Methodology
We assume that we observe $n$ i.i.d observations from $X$ and $Y$ generated by model (1). We concatenate them row-wise into the design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ and the vector of responses $\mathbf{Y}\in\mathbb{R}^{n}$. Classical regression methods ignoring the confounding would estimate $f^{0}$ by minimizing the least squares objective $\hat{f}\coloneqq\text{arg\ min}_{f\in\mathcal{F}}\|\mathbf{Y}-f(\mathbf{X})\|_{2}^{2}$, or a regularized version thereof, for a suitable class $\mathcal{F}$ of functions. This yields an estimate of $\text{arg\ min}_{f\in\mathcal{F}}\mathbb{E}[(Y-f(X))^{2}]=\mathbb{E}[Y|X]=f^{0}(X)+\delta^{T}\mathbb{E}[H|X]$, which is a biased estimate for $f^{0}$ in model (1). We apply a spectral transformation to the least squares objective to remove this confounding bias. Let $Q\in\mathbb{R}^{n\times n}$ be a transformations matrix that depends on the data $\mathbf{X}$. Examples are the trim-transform and the PCA adjustment (Ćevid et al. 2020, Guo et al. 2022, Scheidegger et al. 2025). We use the trim-transform in our empirical results in Section 5, which limits all the singular values of $\mathbf{X}$ to some constant $\tau$, but the methodology can be applied using other spectral transformations. Let $\mathbf{X}=UDV^{T}$ be the singular value decomposition of $\mathbf{X}$, where $U\in\mathbb{R}^{n\times r}$, $D\in\mathbb{R}^{r\times r}$, $V\in\mathbb{R}^{p\times r}$, where $r\coloneqq\min(n,p)$ is the rank of $\mathbf{X}$. The trim transform $Q$ is then defined as
@@
Figure 2: A random realization according to the confounding model (1) with non-linear $f^{0}$ as in (13) and with the same parameter as in Section 5. On the left, we show $f^{0}(\mathbf{X})$ against $\mathbf{Y}$; on the right, the spectrally transformed versions are shown against each other, that is, $Qf^{0}(\mathbf{X})$ versus $Q\mathbf{Y}$. In both visualizations, the line with a slope equal to one, which corresponds to perfect correlation, is shown as a dashed line.
@@ 4 SDForest Algorithm
In principle, our algorithm works similarly to the original CART algorithm (Breiman et al. 2017) and Random Forests (Breiman 2001). The main difference is that we minimize the spectral objective (3) instead of the classical mean squared error. This results in additional challenges. For example, we can no longer decompose the loss and treat leaves independently but must estimate tree levels globally using linear regression.
@@ 4.1 Spectrally Deconfounded Tree
where $(R_{m})_{m=1}^{M}$ are regions dividing the space of $\mathbb{R}^{p}$ into $M$ rectangular parts. Each region has response level $c_{m}\in\mathbb{R}$. We can write the sample version as $f^{0}(\mathbf{X})=\mathcal{P}c$ where $\mathcal{P}\in\{0,1\}^{n\times M}$ is an indicator matrix encoding to which region an observation belongs, i.e. $\mathcal{P}_{i,m}=1$ if the $i$th row of $\mathbf{X}$ belongs to $R_{m}$ and $\mathcal{P}_{i,m}=0$ otherwise. We refer to $\mathcal{P}$ also as a partition, slightly abusing terminology. The vector $c=(c_{1},\ldots,c_{M})\in\mathbb{R}^{M}$ contains the levels corresponding to the different regions. We can estimate $\hat{f}$ in the spirit of (3) with
$(\hat{\mathcal{P}},\hat{c})\coloneqq\text{arg\ min}_{\mathcal{P}^{\prime}\in\{0,1\}^{n\times M},c^{\prime}\in\mathbb{R}^{M}}\frac{\|Q(\mathbf{Y}-\mathcal{P^{\prime}}c^{\prime})\|_{2}^{2}}{n},$ (9)
where $\mathcal{P}^{\prime}$ has to follow a repeated splitting of the space of $\mathbb{R}^{p}$ into rectangular regions. $M$ is fixed here but has to be estimated in practice. $\hat{M}$ also has to be regularized so that it does not go to $n$, and we overfit by having a region for each observation. If the partition $\mathcal{P}$ is known, the spectral objective (3) becomes
@@
where $\mathbf{\tilde{Y}}\coloneqq Q\mathbf{Y}$ and $\mathcal{\tilde{P}}\coloneqq Q\mathcal{P}$. This is a linear regression problem, and we estimate $c$ by
@@
This contrasts the classical CART setting, where, given a partition $\mathcal{P}$, we estimate $c$ with the mean per region $\mu_{m}$. In Equation (10), each $\hat{c}_{m}$ depends not only on the observations belonging to $R_{m}$ but on all the observations due to the rotation and scaling of the spectral transformation. For estimating $\hat{\mathcal{P}}$ we propose Algorithm 1. Algorithm 1 estimates the partition using a tree structure and a repeated splitting of leaves. Since comparing all possibilities for $\mathcal{P}$ is impossible, we grow a tree greedily. Given a current tree, we iterate over all regions and all possible splits. We choose the one that reduces the spectral loss (9) the most, using a subroutine described in Section 4.2, and estimate after each split all the region estimates $\hat{c}$ using Equation (10) in line 17 of Algorithm 1. As the number of leaves grows, the more possible splits we have to compare. In the classical CART setting, an optimal split in a region and its loss decrease remain the same even when another region is divided. Because the spectral transformation $Q$ induces dependency between the observations, the operations for regions depend on all the data. Therefore, to find the next optimal split, we would have to estimate the loss decrease in all potential splits anew in every iteration. Despite this, we reuse the previously estimated loss decrease to save computation time and argue that the change is small and still results in a reasonable split; see also Appendix A (note that we still re-estimate $\hat{c}$ after each iteration). This means that, for each iteration, we only calculate the potential loss decreases for all potential splits in the two new regions. This is repeated until the loss decreases less than a minimum loss decrease after a split. The minimum loss decrease equals a cost-complexity parameter $\mathsf{cp}$ times the initial loss when only an overall mean is estimated. The cost-complexity parameter $\mathsf{cp}$ controls the complexity of a regression tree and acts as a regularization parameter. This should stop $\widehat{M}$ from growing up to $n$. Algorithm 1 results in a regression tree that has been grown aiming to minimize the spectral loss (9) following Equation (10).
4.2 Subroutine
@@
In the $M$th iteration of Algorithm 1, we have the indicator matrix $\hat{\mathcal{P}}=\hat{\mathcal{P}}^{M}\in\{0,1\}^{n\times M}$ with entries $\hat{\mathcal{P}}^{M}_{i,m}=1$ if and only if the $i$th observation lies in region $m$. We encode a candidate split in the region $b$ of the partition using covariate $j$ at the splitting value $s$ by $e\in\{0,1\}^{n}$ where $e$ must be in the support of the $b$th column of $\hat{\mathcal{P}}^{M}$ (i.e. the indices of $1$s in $e$ are a subset of the indices of $1$s of the $b$th column of $\hat{\mathcal{P}}^{M}$)) and the entries depend on $j$ and $s$. A candidate split results then in a candidate partition $\mathcal{P}^{M+1}(e)$ with $M-1$ columns equal to columns $\hat{\mathcal{P}}^{M}$, the $b$th column equal to the $b$th column of $\hat{P}^{M}$ minus $e$ and an additional column equal to $e$. In lines 6-12 of Algorithm 1, we seek to find the optimal $e$ among a large number of candidate splits such that the spectral objective $\|QY-Q\mathcal{P}^{M+1}(e)\hat{c}\|_{2}^{2}$ is minimal, where $\hat{c}$ is the least squares estimator $\hat{c}\coloneqq\arg\min_{c\in\mathbb{R}^{M+1}}\|QY-Q\mathcal{P}^{M+1}(e)c\|_{2}^{2}$. Naively, for every candidate $e$, one would update the indicator matrix $\hat{\mathcal{P}}$, calculate the corresponding least squares estimator $\hat{c}$ and plug it in to obtain the loss decrease. Using the following procedure, the decrease in loss can be calculated more efficiently.
@@
Let $\Pi_{M+1}(e)\in\mathbb{R}^{n\times n}$ be the orthogonal projection on $\textup{span}(Q\mathcal{P}^{M+1}(e))=\textup{span}(Q\hat{\mathcal{P}}^{M},Qe)=\textup{span}(u_{1},\ldots,u_{M},u_{M+1}(e))$. We seek for $e$ that minimizes $\|Q\mathbf{Y}-Q\mathcal{P}^{M+1}(e)\hat{c}\|_{2}^{2}=\|(I_{n}-\Pi_{M+1}(e))Q\mathbf{Y}\|_{2}^{2}$. But because $(u_{1},\ldots,u_{M},u_{M+1}(e))$ is an orthonormal set, it follows that
@@
13:     $(b^{*},j^{*},s^{*})\leftarrow\text{arg\ max }{\alpha_{b,j,s}}$triangleright optimal split over $b\in\{1,\ldots,M\}$
@@
16:     $\hat{\mathcal{P}}^{*}\leftarrow$ splitting $\hat{\mathcal{P}}$ at $(b^{*},j^{*},s^{*})$triangleright resulting in $\hat{\mathcal{P}}^{*}\in\mathbb{R}^{n\times(M+1)}$
@@
24:         $\mathcal{B}\leftarrow(b^{*},M+1)$triangleright the new partitions

### v3

@@ Abstract
We introduce a modification of Random Forests to estimate functions when unobserved confounding variables are present. The technique is tailored for high-dimensional settings with many observed covariates. We employ spectral deconfounding techniques to minimize a deconfounded version of the least squares objective, resulting in the Spectrally Deconfounded Random Forests (SDForests). We demonstrate how the omitted variable bias in estimating a direct effect approaches zero, assuming dense confounding and high-dimensional data. We compare the performance of SDForests with that of classical Random Forests in a simulation study and a semi-synthetic setting using single-cell gene expression data. Empirical results suggest that SDForests outperform classical Random Forests in estimating the direct regression function, even if the theoretical assumptions are not perfectly met, and that SDForests and classical Random Forests have comparable performance in the non-confounded case. We provide an R-Package for SDForest, and supplementary materials for this article are available online.
@@ 1 Introduction
Random Forests (Breiman 2001) and their variations, such as Random Survival Forests (Hothorn 2005, Taylor 2011), Quantile Regression Forests (Meinshausen 2006), or distributional versions of Random Forests (Hothorn & Zeileis 2021, Ćevid et al. 2022) are successfully applied to a wide range of datasets. In many cases of observational data, however, problems with ”omitted variable bias” (Cinelli & Hazlett 2020, Wilms et al. 2021) arise. This means that a bias is induced in estimating relationships using standard Random Forest versions when covariates that correlate with other covariates and the response are not observed and included.
In the setting of causality, this can be viewed as a confounded causal relationship with unobserved confounders (Pearl 2009, Peters et al. 2016). A popular approach to deal with unobserved confounding is to use instrumental variables (IV) regression techniques (Bowden et al. 1990, Angrist et al. 1996, Stock & Trebbi 2003). Finding enough strong and valid instrumental variables can be challenging, especially if many covariates with potential effects on the response are observed, since the number of instruments must be as large as the number of effective covariates.
Another possible way of reducing the hidden confounding bias, which we will adopt here, is to make some kind of ”dense confounding effect” assumption, meaning that the non-observed factors or confounders affect most of the covariates. Then, a standard approach is to estimate the hidden confounding using methods from high-dimensional factor analysis (Bai 2003) and explicitly adjust for them, see for example Leek & Storey (2007), Gagnon-Bartsch & Speed (2012), Fan et al. (2024) for approaches in this direction. Instead of estimating the latent factors explicitly, one can adjust for them implicitly using spectral transformations (Ćevid et al. 2020). Applying such spectral transformations, especially the trim transform, which we introduce later, does not require any tuning and is computationally very fast since it is a simple and explicit function of the singular value decomposition of the design matrix.
@@ 1.1 Our Contribution
Our contribution is a Random Forest algorithm that is able to address, at least partially, the problem of hidden confounding. Our proposal combines the great advantages and flexibility of standard Random Forests with spectral deconfounding techniques for addressing bias from unobserved factors or confounding. The latter was originally proposed for linear models by Ćevid et al. (2020), Guo et al. (2022) and was further developed for nonlinear additive models by Scheidegger et al. (2025). The application of spectral deconfounding techniques for Random Forests is novel.
We develop a new algorithm with the R-package SDModels (Ulmer & Scheidegger 2025) and show its performance in estimating the direct and unconfounded relationship between the observed covariates and the response. We demonstrate that in the presence of hidden confounding, our method, the Spectrally Deconfounded Random Forest (SDForest), outperforms the standard Random Forests in many aspects. If no latent factor or confounding exists, our SDForests and standard Random Forests perform similarly, perhaps with a minimal edge in favor of the classical algorithm. Thus, if one is unsure whether hidden confounding is present, there is much to gain but not much to lose.
@@ 2 Confounding Model
Here, $X\in\mathbb{R}^{p}$ are the predictors, $Y\in\mathbb{R}$ is the response, and $H\in\mathbb{R}^{q}$ are unobserved hidden confounding factors. We assume that the confounder influences $X$ with a linear effect $\Gamma\in\mathbb{R}^{q\times p}$ and $Y$ with a linear effect $\delta\in\mathbb{R}^{q}$ (see Appendix C for a note on non-linear confounding); without loss of generality, we can assume $\mathrm{Cov}(H)=I$. Furthermore, $\nu$ is a random variable with mean zero and variance $\sigma_{\nu}^{2}$, and $E$ is a random vector with mean zero and covariance $\Sigma_{E}$, and $E$ and $H$ are uncorrelated. The error term $E$ can be viewed as the unconfounded predictor if $\Gamma$ equals zero. Finally, $f^{0}\in\mathcal{F}$, where $\mathcal{F}$ is some class of functions from $\mathbb{R}^{p}$ to $\mathbb{R}$. The function $f^{0}$ encodes the direct causal relationship of interest, describing the causal relation of $X$ on $Y$. The described model is visualized as a graph in Figure 1.
[…]
@@ 3 Generic Methodology
We assume that we observe $n$ i.i.d. observations from $X$ and $Y$ generated by model (1). We concatenate them row-wise into the design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ and the vector of responses $\mathbf{Y}\in\mathbb{R}^{n}$. Classical regression methods ignoring the confounding would estimate $f^{0}$ by minimizing the least squares objective $\hat{f}\coloneqq\text{arg\ min}_{f\in\mathcal{F}}\|\mathbf{Y}-f(\mathbf{X})\|_{2}^{2}$, or a regularized version thereof, for a suitable class $\mathcal{F}$ of functions. This yields an estimate of $\text{arg\ min}_{f\in\mathcal{F}}\mathbb{E}[(Y-f(X))^{2}]=\mathbb{E}[Y|X]=f^{0}(X)+\delta^{T}\mathbb{E}[H|X]$, which is a biased estimate for $f^{0}$ in model (1). We apply a spectral transformation to the least squares objective to remove this confounding bias. Let $Q\in\mathbb{R}^{n\times n}$ be a transformation matrix that depends on the data $\mathbf{X}$. Examples are the trim-transform and the PCA adjustment (Ćevid et al. 2020, Guo et al. 2022, Scheidegger et al. 2025). We use the trim-transform in our empirical results in Section 5, which limits all the singular values of $\mathbf{X}$ to some constant $\tau$, but the methodology can be applied using other spectral transformations. Let $\mathbf{X}=UDV^{T}$ be the singular value decomposition of $\mathbf{X}$, where $U\in\mathbb{R}^{n\times r}$, $D\in\mathbb{R}^{r\times r}$, $V\in\mathbb{R}^{p\times r}$, where $r\coloneqq\min(n,p)$ is the rank of $\mathbf{X}$. The trim transform $Q$ is then defined as
@@
Figure 2: A random realization according to the confounding model (1) with non-linear $f^{0}$ as in (13) and with the same parameter as in Section 5. On the left, we show $f^{0}(\mathbf{X})$ against $\mathbf{Y}$; on the right, the spectrally transformed versions are shown against each other, that is, $Qf^{0}(\mathbf{X})$ versus $Q\mathbf{Y}$. In both visualizations, the line with a slope of one, corresponding to equality, is shown as a dashed line.
@@ 3.1 Assumptions and Technical Motivation
Assume the confounding model (1) and assume that the conditions (4), (5), (6), and (7) hold. Then, it holds that
@@
Theorem 1 justifies the minimization of the spectral objective (3) to remove the confounding bias. Additionally, to effectively estimate $f^{0}(\cdot)$ in the high-dimensional setting, it seems crucial that $f^{0}(\cdot)$ is sparse. In fact, for the theory in the linear and the additive case (Ćevid et al. 2020, Guo et al. 2022, Scheidegger et al. 2025), sparsity is a central requirement for consistency of spectral deconfounding, so we expect sparsity to also play an important role here.
@@ 4 SDForest Algorithm
In principle, our algorithm works similarly to the original CART algorithm (Breiman et al. 2017) and Random Forests (Breiman 2001). CART minimizes the mean squared error by greedily dividing the space $\mathbb{R}^{p}$ into rectangular parts. Each region then has a response level resulting in a function as in Equation (8) below. CART starts with a single region and then searches in all the variables and along their support for the split that minimizes the mean squared error. This process is repeated for both subsequent regions and continues in the same manner. The mean squared error has the nice property that the loss decomposes, meaning that for the next optimal split in a region, CART only needs to look at the samples belonging to this region. The main difference between our algorithm and CART is that we minimize the spectral objective (3) instead of the classical mean squared error. This results in additional challenges as the spectral objective is no longer decomposable.
@@ 4.1 Spectrally Deconfounded Tree
where $(R_{m})_{m=1}^{M}$ are regions dividing the space of $\mathbb{R}^{p}$ into $M$ disjoint rectangular parts. Each region has a response level $c_{m}\in\mathbb{R}$. We can write the sample version as $f^{0}(\mathbf{X})=\mathcal{P}c$ where $\mathcal{P}\in\{0,1\}^{n\times M}$ is an indicator matrix encoding to which region an observation belongs, i.e. $\mathcal{P}_{i,m}=1$ if the $i$th row of $\mathbf{X}$ belongs to $R_{m}$ and $\mathcal{P}_{i,m}=0$ otherwise. We refer to $\mathcal{P}$ also as a partition, slightly abusing terminology. The vector $c=(c_{1},\ldots,c_{M})\in\mathbb{R}^{M}$ contains the levels corresponding to the different regions. We can estimate $\hat{f}$ in the spirit of (3) with
[…]
where $\mathfrak{P}\in\{0,1\}^{n\times M}$ and $\mathfrak{P}$ has to result from a repeated splitting of the space of $\mathbb{R}^{p}$ into rectangular regions. Repeated splitting means that $\mathfrak{P}$ can be represented by a tree structure. Each branching in the tree corresponds to the splitting of $\mathbb{R}^{p}$ at a variable $j$ and a split point $s$.
$M$ is fixed here, but has to be estimated in practice. The estimator $\hat{M}$ for $M$ must be regularized to prevent overfitting. If the partition $\mathcal{P}$ is known, the spectral objective (3) becomes
@@
where $\mathbf{\tilde{Y}}\coloneqq Q\mathbf{Y}$ and $\mathcal{\tilde{P}}\coloneqq Q\mathcal{P}$. This is a linear regression problem, and we compute $\hat{c}$ with
@@
In the spirit of CART, we propose Algorithm 1 to find a tree representing $\hat{\mathcal{P}}$ and $\hat{c}$. While growing the tree, we try to find the next best split. The next best split needs to be chosen among all the current leaves, all variables, and somewhere in $\mathbb{R}$. To calculate how much each split reduces the spectral loss (3), we employ a subroutine described in the next section. Using these loss decreases, we select the split that yields the maximum reduction in training loss.
In a subsequent step, the splits in the unused leaves resulting in the maximal loss decrease may no longer be the same. This is due to the induced dependency by the spectral transformation. However, to save computation time, we do not recalculate the optimal split and loss decrease for all leaves, but instead calculate optimal splits and loss decreases only for newly created leaves. For the old leaves, we reuse the previously calculated ones as approximations. We argue that the change is minor and still yields reasonable splits. (In the software, both options are available.) Appendix A shows a comparison of the performance of this approximation. Note that we still re-estimate $\hat{c}$ after each iteration.
To avoid overfitting, a split is only carried out if it reduces the spectral loss (3) enough. The strength of this regularization is controlled by the cost-complexity parameter $\mathsf{cp}$. Only splits that reduce the loss at least as much as $\mathsf{cp}$ times the initial loss, i.e., $\mathsf{cp}\times\frac{\|Q(\mathbf{Y}-\bar{\mathbf{Y}})\|_{2}^{2}}{n}$.
4.2 Subroutine to Evaluate a Split
@@
This section explains how we can search for and evaluate potential splits more efficiently. In the $M$th iteration of Algorithm 1, we have the indicator matrix $\hat{\mathcal{P}}=\hat{\mathcal{P}}^{M}\in\{0,1\}^{n\times M}$ with entries $\hat{\mathcal{P}}^{M}_{i,m}=1$ if and only if the $i$th observation lies in region $m$. We encode a candidate split in the region $b$ of the partition using covariate $j$ at the splitting point $s$ by $e\in\{0,1\}^{n}$ where $e$ must be in the support of the $b$th column of $\hat{\mathcal{P}}^{M}$ (i.e. the indices of $1$s in $e$ are a subset of the indices of $1$s of the $b$th column of $\hat{\mathcal{P}}^{M}$)) and the entries depend on $j$ and $s$. A candidate split results in a candidate partition $\mathcal{P}^{M+1}(e)$ with M1M-1 columns equal to columns $\hat{\mathcal{P}}^{M}$, the $b$th column equal to the $b$th column of $\hat{P}^{M}$ minus $e$, and an additional column equal to $e$.
In lines 6 to 12 of Algorithm 1, we seek to find the optimal $e$ among a large number of candidate splits such that the spectral objective $\|QY-Q\mathcal{P}^{M+1}(e)\hat{c}\|_{2}^{2}$ is minimal, where $\hat{c}$ is the least squares estimator $\hat{c}\coloneqq\arg\min_{c\in\mathbb{R}^{M+1}}\|QY-Q\mathcal{P}^{M+1}(e)c\|_{2}^{2}$. Naively, for every candidate $e$, one would update the indicator matrix $\hat{\mathcal{P}}$, calculate the corresponding least squares estimator $\hat{c}$ and plug it in to obtain the loss decrease. Using the following procedure, the decrease in loss can be calculated more efficiently.
@@
Let $\Pi_{M+1}(e)\in\mathbb{R}^{n\times n}$ be the orthogonal projection on $\textup{span}(Q\mathcal{P}^{M+1}(e))=\textup{span}(Q\hat{\mathcal{P}}^{M},Qe)=\textup{span}(u_{1},\ldots,u_{M},u_{M+1}(e))$. We seek for $e$ that minimizes $\|Q\mathbf{Y}-Q\mathcal{P}^{M+1}(e)\hat{c}\|_{2}^{2}=\|(I_{n}-\Pi_{M+1}(e))Q\mathbf{Y}\|_{2}^{2}$. Because $(u_{1},\ldots,u_{M},u_{M+1}(e))$ is an orthonormal set, it follows that
@@
13:  $(b^{*},j^{*},s^{*})\leftarrow\text{arg\ max }{\alpha_{b,j,s}}$ $\triangleright$ optimal split over $b\in\{1,\ldots,M\}$
@@
16:  $\hat{\mathcal{P}}^{*}\leftarrow$ splitting $\hat{\mathcal{P}}$ at $(b^{*},j^{*},s^{*})$ $\triangleright$ resulting in $\hat{\mathcal{P}}^{*}\in\mathbb{R}^{n\times(M+1)}$
@@
24:   $\mathcal{B}\leftarrow(b^{*},M+1)$ $\triangleright$ the new partitions